You have a solution mixed at 8 g/ml and need 81 gallons mixed at 4 g/ml. How many gallons of water need to be added to the 8 g/ml solution? Round answers to the nearest tenth of a gallon
Answer by richwmiller(17219)

from view of solution
81*0.08=0.04*(81+x)
6.48=3.24+0.04*x
6.48-3.24=0.04*x
3.24=0.04*x
x=81
check
81*0.08=0.04*(81+x)
81*0.08=0.04*(162)
6.48=6.48
ok
or from view of water
1*y+0.92*81=0.96(81+y)
1*y+74.52=77.76+0.96y
1*y-0.96y=77.76-74.52
0.04y=3.24
y=3.24/0.04
y=81